Three years ago, Joe's age was 1 year more than twice Karen's. Six years from now, his age will be 10 years more than half her age. how old is Joe?
Answer by stanbon(75887)

Three years ago, Joe's age was 1 year more than twice Karen's. Six years from now, his age will be 10 years more than half her age. how old is Joe?
-------
Equations:
(j-3) = 2(k-3)+1
(j+6) = (1/2)(k+6)+10
-----
Rearrange the 1st equation:
j-3 = 2k-6+1
j-3 = 2k-5
j = 2k-2
---
Substitute into the 2nd equation and solve for "k":
(2k-2+6) = (1/2)(k + 6) + 10
Multiply thru by 2:
2(2k+4) = k+6+20
4k+8 = k+26
3k = 18
k = 6 (Karen's age now)
---
Solve for "j":
j = 2k-2
j = 12-2
j = 10 (Joe's age now)
===========================
Cheers,
Stan H.
===============